OpenStudy (anonymous):

can someone explain to me why (2z+5)/2(z+5) simplifies to 1/2?

OpenStudy (anonymous):

Are you working with limits

OpenStudy (anonymous):

Like non-permissible values? yes. the original equation was: \[\frac{ 4z^2-25 }{ 2z^2-13z+20 }\times \frac{ z-4 }{ 4z+10 }\] I simplified it to: \[\frac{ 2z+5 }{ 2(z+5) }\] but the answer says it's 1/2 and i'm not sure how to get there?

OpenStudy (anonymous):

@jdoe0001 can probably help you. lol

OpenStudy (jdoe0001):

you can't

OpenStudy (jdoe0001):

hmm.... maybe there's more context.... thus it comes down to 1/2... but \(\bf \cfrac{2z+5}{2(z+5)}\) doesn't simplify further as far as I can tell

OpenStudy (jdoe0001):

as OrionsBelt said... if it's say \(\bf lim_{z\to 0}\) then yes, you'd end up with \(\bf \cfrac{1}{2}\)

OpenStudy (anonymous):

thanks